If there is a square of side 'a'. At the centre of the square there is situated a tower of height 'h', from the top of the tower the angle at the top and the line joining the vertex is 60 degrees. Find the height of the tower?
1
Expert's answer
2011-06-09T11:59:32-0400
AC = √(AD2 + DC2) = √(AD2) = a√2
OC = 1/2 AC = a/√2
OH = OC/tan(60) = a/√(2*3) = a/√6